OpenStudy (anonymous):

What is the exact value of sqrt of 75/36?

OpenStudy (anonymous):



OpenStudy (anonymous):

you are awesome :)

jhonyy9 (jhonyy9):

5sqrt3 ----- 6

jhonyy9 (jhonyy9):

because 75 is 25*3 =5*5*3

OpenStudy (anonymous):

You can simplify this to \[\frac{5\sqrt{3}}{\sqrt{36}}\] Which is \[\frac{5\sqrt{3}}{6}\] or about 52

jhonyy9 (jhonyy9):

36=6*6

OpenStudy (anonymous):

you guys are great :) Thank you all alot! <3